### v1

@@ Abstract
The neural ordinary differential equation (ODE) framework has shown considerable promise in recent years in developing highly accelerated surrogate models for complex physical systems characterized by partial differential equations (PDEs). For PDE-based systems, state-of-the-art neural ODE strategies leverage a two-step procedure to achieve this acceleration: a nonlinear dimensionality reduction step provided by an autoencoder, and a time integration step provided by a neural-network based model for the resultant latent space dynamics (the neural ODE). This work explores the applicability of such autoencoder-based neural ODE strategies for PDEs in which advection terms play a critical role. More specifically, alongside predictive demonstrations, physical insight into the sources of model acceleration (i.e., how the neural ODE achieves its acceleration) is the scope of the current study. Such investigations are performed by quantifying the effects of both autoencoder and neural ODE components on latent system time-scales using eigenvalue analysis of dynamical system Jacobians. To this end, the sensitivity of various critical training parameters – de-coupled versus end-to-end training, latent space dimensionality, and the role of training trajectory length, for example – to both model accuracy and the discovered latent system timescales is quantified. This work specifically uncovers the key role played by the training trajectory length (the number of rollout steps in the loss function during training) on the latent system timescales: larger trajectory lengths correlate with an increase in limiting neural ODE time-scales, and optimal neural ODEs are found to recover the largest time-scales of the full-order (ground-truth) system. Demonstrations are performed across two fundamentally different unsteady fluid dynamics configurations influenced by advection: (1) the Kuramoto-Sivashinsky equations, and (2) Hydrogen-Air channel detonations (the compressible reacting Navier-Stokes equations with detailed chemistry).
journal:
@@ 1 Introduction
Numerical solutions of partial differential equations (PDEs), if available, can be used by domain scientists to not only probe complex physical behavior at unprecedented levels of accuracy and detail, but also to accelerate design optimization workflows for engineering devices. Simulating fluid dynamics, for example, requires numerically solving the Navier-Stokes PDEs moin_dns . At industrial operating conditions of interest – those that characterize flows over aircraft wings, in scramjets, and in gas turbine combustion chambers, for instance – these equations admit multi-scale and multi-physics behavior stemming from interactions between turbulence, shock waves, and potential chemical reactions. As a result, to generate reliable simulations of such devices, PDE solution procedures need to resolve all length-scales and time-scales that characterize these physical phenomena. These spatiotemporal resolution requirements render (a) long-time direct numerical simulation (DNS) of the aforementioned systems infeasible, and (b) real-time simulation-based actuation or control strategies impractical, despite recent advances in supercomputing technology and physics simulation hardware venkat_emergingtrends ; exascale .
@@
Reduced-order models (ROMs) are a class of modeling approaches that seek to eliminate intrinsic costs associated with physics-based simulations lucia2004reduced , with the goal of enabling long-time simulation capability. The general ROM objective is to achieve drastic reduction in the PDE-derived dynamical system, which is the high-dimensional nonlinear set of ordinary differential equations (ODEs) produced from PDE discretization. In the context of fluid flows, this dynamical system describes the evolution of a so-called state vector described by turbulent fluid density, momentum, and species concentration fields on a grid, and can readily reach on the order of hundreds of millions (and even billions) of degrees-of-freedom. As such, to achieve model-based reduction, ROMs typically leverage a two-step approach: an offline projection step which generates the reduced representation of the state vector, and an online forecasting step which generates the time-evolution of the reduced state. Within this scope, both physics-based and data-based ROMs can be constructed.
@@
Physics-based ROMs for fluid flows typically leverage linear projection operations to resolve only the large scales, which are assumed to contain a majority of the system energy; the effect of the unresolved (small) scales on the resolved dynamics is then modeled. A classic example is large-eddy simulation (LES) fox_les ; fureby_les , where the projection operation (the mechanism for dimensionality reduction) is a non-invertible spatial filter ideal_les , and the forecasting step requires solving a filtered version of the Navier-Stokes equations. The LES closure model – which captures the effects of the unresolved (small) scales on the resolved dynamics – can come from either phenomenological algebraic relations dynamic_model or statistical closures ideal_les ; adrian_1988 . Other physics-based ROMs achieve dimensionality reduction in different ways, such as through inertial manifold assumptions foias_aim ; maryam_aim , Koopman operator theory parish_2017 , and more intrusive alterations to the governing PDEs (e.g., flamelet models used in combustion modeling pitsch_arfm and two-dimensional turbulence used in climate modeling 2d_turb ).
@@
Data-based ROMs, on the other hand, rely on samples of the full-dimensional state vector (i.e., fluid flow snapshots) to produce projection operators in an optimization (or training) step. Although this incurs a large offline computational cost not present in physics-based counterparts, these methods have been shown in recent years to produce significantly larger levels of dimensionality reduction, such that this training cost can be offset by immense levels of speedup achieved during the forecasting stage kutz2017deep ; karthik_arfm ; venkat_emergingtrends . In the context of data-based ROMs, the reduced space is termed the latent space. The initial projection step produces the (reduced) latent variables, and the forecasting step requires modeling the dynamics of these latent variables.
@@
Data-based ROMs have a history spanning several decades, and their ability to capture complex physics contained in training datasets has led to their increased adoption. Methods rooted in modal decomposition – including proper orthogonal decomposition (POD) berkooz_pod , dynamic mode decomposition (DMD) schmid_dmd_arfm , resolvent analysis towne2018spectral , and cluster-based methods gunzburger_cvt , among others – derive basis functions directly from data, which translates to latent spaces generated by linear projection operations. The properties of the resultant latent variables naturally vary based on the method used to produce the basis. In POD, a basis is produced from eigenvectors of the covariance matrix of the data, resulting in latent variables that are disentangled and optimally preserve the system energy berkooz_pod . In DMD and resolvent analysis, the linear projection generates latent variables described by characteristic frequencies, similar to traditional Fourier-based methods schmid_dmd_arfm . Cluster-based methods use data partitioning strategies to produce latent variables that are symbolic encodings of the system state kaiser_crom ; shivam_timeaxis . The evolution of latent variables in all of these methods can be modeled in a physics-derived manner through Galerkin projection of the basis onto the underlying PDEs pod_galerkin , or in a data-based manner through the utilization of machine learning methods romit_spod .
@@
All of the methods discussed above leverage linear projection operations to achieve reduction. These methods have been successfully used to produce ROMs for diffusion-dominated problems, such as turbulent flows in canonical configurations and simpler PDEs without advection terms, but face difficulites for advection-dominated problems (e.g., high Reynolds number turbulent flows and shock-containing flows) characterized by slow decaying Kolmogorov n-width ohlberger2015reduced ; LEE2020108973 .
@@
The utilization of data-based ROMs relying on nonlinear projection operations have been shown to overcome these limitations, extending both compression and forecasting capabilities of linear approaches romit_cnn_advection ; lee2020model ; glaws2020deep . The backbone of these methods is the autoencoder, a compression approach that leverages the expressive power of neural networks to generate robust latent space representations autoencoder_review . Autoencoders rely on two components: the encoder, which serves as the nonlinear projection that moves the high-dimensional state into its latent representation, and the decoder, which seeks to undo the action of the encoder by recovering the full state from the latent variables. Due to the generalizable nature of neural networks, autoencoders can take many forms tailored to the application at hand: for example, architectures can leverage multi-layer perceptrons bethany_koopman , convolutional neural networks romit_cnn_advection ; lee2020model , graph neural networks gunzburger_gnn ; shivam_gnn , and transformers liu2019transformer .
@@
The success of neural network based autoencoders has driven efforts to create purely data-based predictive models in latent space that offer unprecedented levels of speedup over physics-based counterparts lee_parish_2020 . A core advantage is that such models can leverage real-world observations (e.g., from experimental diagnostics and operational sensor streams), which is critical for applications that are too expensive to simulate directly or do not have a solidified set of governing equations  xu2020multi ; yu2019non ; fresca2021comprehensive ; duraisamy2015new ; hesthaven2018non . The goal of these forecasting models is to operate in concert with the autoencoder by using data to learn the dynamics of the latent variables. The latent dynamical systems can be modeled using nonlinear romit_lstm ; reddy_rnn or linear bethany_koopman ; kaiser_crom surrogate forecasting models; such models have been used to accelerate advection dominated fluid flow romit_cnn_advection , chaotic dynamical systems bethany_koopman , and stiff chemical kinetic simulations ji2021stiff . These methods, however – e.g., those based on strategies like recurrent neural networks (RNNs) romit_cnn_advection , residual networks (ResNets) jiang2021deep , and latent Koopman representations yeung2019learning – typically learn the latent dynamics in the context of discrete and explicit temporal discretization, which can limit predictive capability.
@@
The goal of this study is to develop surrogate models for advection-dominated systems in the neural ODE framework chen2018neural , which offers unique advantages over the above mentioned approaches. More specifically, the objective here is twofold: (1) to develop a neural ODE based latent dynamics model for advection-dominated problems in a purely data-based setting, and (2) to conduct a detailed analysis of the discovered dynamics in latent space using timescale characterizations. Before outlining the specific contributions of this work, the neural ODE strategy as it relates to these objectives is first summarized.
@@
Since its introduction in Ref. chen2018neural , the neural ODE strategy has been cemented as a powerful scientific machine learning tool to model the evolution of dynamical systems using neural networks. Instead of directly enforcing discrete temporal representations (e.g., as used in residual networks jiang2021deep or recurrent neural networks reddy_rnn ), neural ODEs learn a continuous representation of nonlinear system dynamics. In other words, the instantaneous right-hand-side is modeled as a nonlinear function via a neural network, allowing the framework to leverage existing, vetted time-integration schemes to execute the forecasting step (the time integration scheme is separated from the dynamics model) chemnode ; Finlay2020 ; inproceedings ; kumar2023posteriori . With autoencoder-provided latent spaces, the neural ODE conveniently outputs a functional form for the instantaneous rate-of-change of the latent variables, which is useful for reduced-order modeling applications. This autoencoder-based neural ODE strategy been used to develop accelerated surrogate models in a variety of physical applications, including chaotic fluid flows Dutta2021 , advection-dominated PDEs Linot2023 ; Wan2023 , and stiff chemical kinetics dikeman2022stiffness ; Vijay23 .
@@
Ultimately, previous work has shown how the combination of autoencoders with neural ODEs can be used to generate highly accelerated reduced-order models of physical systems. Despite this, the source of acceleration in the overall modeling strategy remains unclear: the goal of achieving model accuracy from both forecasting and autoencoding perspectives often overshadows the need to identify the contribution of each of these components to the empirically observed model acceleration. Insights into the sources of acceleration provided by the neural ODE based ROM can lead to valuable physical and model-oriented insights, and is the scope of the current study. Recent work has observed the effect of smoothed latent trajectories for advection-dominated systems produced by neural ODE simulations, pointing to a relationship between model acceleration and instrinsic timescale elimination Wan2023 . Similar trends have been shown for neural ODE based surrogate models for stiff chemical kinetics dikeman2022stiffness ; Vijay23 . As such, a more rigorous and quantitative analysis of timescale elimination produced by both autoencoder and neural ODE components is warranted. Additionally, the effect of critical neural ODE training parameters – such as the overall integration time used to evaluate model errors – on the accuracy and degree of timescale elimination produced in the latent space has been largely unexplored in the literature. Lastly, although previous work has demonstrated application of neural ODE based surrogate models for simplified advection-dominated PDEs (e.g., Burger’s equation), extension of this strategy to more complex shock-containing flows remains sparsely explored. To this end, the main contributions of this work are as follows:
@@
    Extend the proposed framework to a challenging and highly advection-dominated test case, specifically 1D detonation wave propagation (considering stiff chemical kinetics), to validate and affirm the observed trends.
@@ 2.1 Surrogate Modeling Task
The ODE in Eq. 2 is interpreted here as a ”ground-truth” representation of the continuous PDE counterpart, meaning the grid is assumed to be resolved enough to properly capture the contribution of all spatiotemporal scales in the physical operators, resulting in a high-dimensional state vector ${\bf u}$ that corresponds to a physical space representation of the state. The motivation for surrogate modeling is that the ground-truth ODE in Eq. 2 is computationally expensive to solve for realistic applications that require fast (near real-time) flow field predictions, such as full-scale design optimization Amsallem2009 ; lieu2007adaptation ; lieu2006reduced or model-predictive control mathelin2009robust ; hovland2008explicit .
@@
It should be noted that accelerated evaluations over the full system using a combined autoencoder and neural ODE strategy can be achieved through both a reduction in system dimensionality through the latent variables (the surrogate ODE operates on a lower-dimensional representation, which facilitates faster evaluations of instantaneous rates), and also an increase in minimum system timescales in the latent space inherent to the functional form of ${\bf G}$ Vijay23 (larger limiting timescales in ${\bf G}$ imply elimination of prohibitive timescales in ${\bf F}$, which in turn allows for larger time steps to be utilized in the simulation procedure). More specifically, previous work has empirically observed how employing a convolutional autoencoder yields smooth trajectories in the latent space for chaotic PDE-based systems Wan2023 . To this end, within the scope of data-based surrogate modeling, the primary emphasis of this work is to (a) provide a rigorous quantitative analysis on how timescale elimination is achieved in the latent space, (b) identify the role played by key neural ODE training parameters on this timescale elimination, and (c) understand the manner in which coupling between autoencoder and neural ODE components contributes to forecasting accuracy and latent timescale sensitivity.

### v3

@@ Abstract
The neural ordinary differential equation (ODE) framework has shown considerable promise in recent years in developing highly accelerated surrogate models for complex physical systems characterized by partial differential equations (PDEs). For PDE-based systems, state-of-the-art neural ODE strategies leverage a two-step procedure to achieve this acceleration: a nonlinear dimensionality reduction step provided by an autoencoder, and a time integration step provided by a neural-network based model for the resultant latent space dynamics (the neural ODE). This work explores the applicability of such autoencoder-based neural ODE strategies for PDEs in which advection terms play a critical role. More specifically, alongside predictive demonstrations, physical insight into the sources of model acceleration (i.e., how the neural ODE achieves its acceleration) is the scope of the current study. Such investigations are performed by quantifying the effects of both autoencoder and neural ODE components on latent system time-scales using eigenvalue analysis of dynamical system Jacobians. To this end, the sensitivity of various critical training parameters – de-coupled versus end-to-end training, latent space dimensionality, and the role of training trajectory length, for example – to both model accuracy and the discovered latent system timescales is quantified. This work specifically uncovers the key role played by the training trajectory length (the number of rollout steps in the loss function during training) on the latent system timescales: larger trajectory lengths correlate with an increase in limiting neural ODE time-scales, and optimal neural ODEs are found to recover the largest time-scales of the full-order (ground-truth) system. Demonstrations are performed across fundamentally different unsteady fluid dynamics configurations influenced by advection: (1) the Kuramoto-Sivashinsky equations (2) Hydrogen-Air channel detonations (the compressible reacting Navier-Stokes equations with detailed chemistry), and (3) 2D Atmospheric flow.
@@ 1 Introduction
Numerical solutions of partial differential equations (PDEs), if available, can be used by domain scientists to not only probe complex physical behavior at unprecedented levels of accuracy and detail, but also to accelerate design optimization workflows for engineering devices. Simulating fluid dynamics, for example, requires numerically solving the Navier-Stokes PDEs [1]. At industrial operating conditions of interest – those that characterize flows over aircraft wings, in scramjets, and in gas turbine combustion chambers, for instance – these equations admit multi-scale and multi-physics behavior stemming from interactions between turbulence, shock waves, and potential chemical reactions. As a result, to generate reliable simulations of such devices, PDE solution procedures need to resolve all length-scales and time-scales that characterize these physical phenomena. These spatiotemporal resolution requirements render (a) long-time direct numerical simulation (DNS) of the aforementioned systems infeasible, and (b) real-time simulation-based actuation or control strategies impractical, despite recent advances in supercomputing technology and physics simulation hardware [2, 3].
@@
Reduced-order models (ROMs) are a class of modeling approaches that seek to eliminate intrinsic costs associated with physics-based simulations [4], with the goal of enabling long-time simulation capability. The general ROM objective is to achieve drastic reduction in the PDE-derived dynamical system, which is the high-dimensional nonlinear set of ordinary differential equations (ODEs) produced from PDE discretization. In the context of fluid flows, this dynamical system describes the evolution of a so-called state vector described by turbulent fluid density, momentum, and species concentration fields on a grid, and can readily reach on the order of hundreds of millions (and even billions) of degrees-of-freedom. As such, to achieve model-based reduction, ROMs typically leverage a two-step approach: an offline projection step which generates the reduced representation of the state vector, and an online forecasting step which generates the time-evolution of the reduced state. Within this scope, both physics-based and data-based ROMs can be constructed.
@@
Physics-based ROMs for fluid flows typically leverage linear projection operations to resolve only the large scales, which are assumed to contain a majority of the system energy; the effect of the unresolved (small) scales on the resolved dynamics is then modeled. A classic example is large-eddy simulation (LES) [5, 6], where the projection operation (the mechanism for dimensionality reduction) is a non-invertible spatial filter [7], and the forecasting step requires solving a filtered version of the Navier-Stokes equations. The LES closure model – which captures the effects of the unresolved (small) scales on the resolved dynamics – can come from either phenomenological algebraic relations [8] or statistical closures [7, 9]. Other physics-based ROMs achieve dimensionality reduction in different ways, such as through inertial manifold assumptions [10, 11], Koopman operator theory [12], and more intrusive alterations to the governing PDEs (e.g., flamelet models used in combustion modeling [13] and two-dimensional turbulence used in climate modeling [14]).
@@
Data-based ROMs, on the other hand, rely on samples of the full-dimensional state vector (i.e., fluid flow snapshots) to produce projection operators in an optimization (or training) step. Although this incurs a large offline computational cost not present in physics-based counterparts, these methods have been shown in recent years to produce significantly larger levels of dimensionality reduction, such that this training cost can be offset by immense levels of speedup achieved during the forecasting stage [15, 16, 2]. In the context of data-based ROMs, the reduced space is termed the latent space. The initial projection step produces the (reduced) latent variables, and the forecasting step requires modeling the dynamics of these latent variables.
@@
Data-based ROMs have a history spanning several decades, and their ability to capture complex physics contained in training datasets has led to their increased adoption. Methods rooted in modal decomposition – including proper orthogonal decomposition (POD) [17], dynamic mode decomposition (DMD) [18], resolvent analysis [19], and cluster-based methods [20], among others – derive basis functions directly from data, which translates to latent spaces generated by linear projection operations. The properties of the resultant latent variables naturally vary based on the method used to produce the basis. In POD, a basis is produced from eigenvectors of the covariance matrix of the data, resulting in latent variables that are disentangled and optimally preserve the system energy [17]. In DMD and resolvent analysis, the linear projection generates latent variables described by characteristic frequencies, similar to traditional Fourier-based methods [18]. Cluster-based methods use data partitioning strategies to produce latent variables that are symbolic encodings of the system state [21, 22]. The evolution of latent variables in all of these methods can be modeled in a physics-derived manner through Galerkin projection of the basis onto the underlying PDEs [23], or in a data-based manner through the utilization of machine learning methods [24].
@@
All of the methods discussed above leverage linear projection operations to achieve reduction. These methods have been successfully used to produce ROMs for diffusion-dominated problems, such as turbulent flows in canonical configurations and simpler PDEs without advection terms, but face difficulites for advection-dominated problems (e.g., high Reynolds number turbulent flows and shock-containing flows) characterized by slow decaying Kolmogorov n-width [25, 26].
@@
The utilization of data-based ROMs relying on nonlinear projection operations have been shown to overcome these limitations, extending both compression and forecasting capabilities of linear approaches [27, 28, 29]. The backbone of these methods is the autoencoder, a compression approach that leverages the expressive power of neural networks to generate robust latent space representations [30]. Autoencoders rely on two components: the encoder, which serves as the nonlinear projection that moves the high-dimensional state into its latent representation, and the decoder, which seeks to undo the action of the encoder by recovering the full state from the latent variables. Due to the generalizable nature of neural networks, autoencoders can take many forms tailored to the application at hand: for example, architectures can leverage multi-layer perceptrons [31], convolutional neural networks [27, 28], graph neural networks [32, 33], and transformers [34].
@@
The success of neural network based autoencoders has driven efforts to create purely data-based predictive models in latent space that offer unprecedented levels of speedup over physics-based counterparts [35]. A core advantage is that such models can leverage real-world observations (e.g., from experimental diagnostics and operational sensor streams), which is critical for applications that are too expensive to simulate directly or do not have a solidified set of governing equations  [36, 27, 37, 38, 39, 40]. The goal of these forecasting models is to operate in concert with the autoencoder by using data to learn the dynamics of the latent variables. The latent dynamical systems can be modeled using nonlinear [41, 42] or linear [31, 21] surrogate forecasting models; such models have been used to accelerate advection dominated fluid flow [27], chaotic dynamical systems [31], and stiff chemical kinetic simulations [43]. These methods, however – e.g., those based on strategies like recurrent neural networks (RNNs) [27], residual networks (ResNets) [44], and latent Koopman representations [45] – typically learn the latent dynamics in the context of discrete and explicit temporal discretization, which can limit predictive capability.
@@
The goal of this study is to develop surrogate models for advection-dominated systems in the neural ODE framework [46], which offers unique advantages over the above mentioned approaches. More specifically, the objective here is twofold: (1) to develop a neural ODE based latent dynamics model for advection-dominated problems in a purely data-based setting, and (2) to conduct a detailed analysis of the discovered dynamics in latent space using timescale characterizations. Before outlining the specific contributions of this work, the neural ODE strategy as it relates to these objectives is first summarized.
@@
Since its introduction in Ref. [46], the neural ODE strategy has been cemented as a powerful scientific machine learning tool to model the evolution of dynamical systems using neural networks. Instead of directly enforcing discrete temporal representations (e.g., as used in residual networks [44] or recurrent neural networks [42]), neural ODEs learn a continuous representation of nonlinear system dynamics. In other words, the instantaneous right-hand-side is modeled as a nonlinear function via a neural network, allowing the framework to leverage existing, vetted time-integration schemes to execute the forecasting step (the time integration scheme is separated from the dynamics model) [47, 48, 49, 50, 51]. A prior version of neural ODEs, integrating feedforward neural networks with numerical methods for dynamical system identification, used a Runge–Kutta neural network framework [52]. With autoencoder-provided latent spaces, the neural ODE conveniently outputs a functional form for the instantaneous rate-of-change of the latent variables, which is useful for reduced-order modeling applications. This autoencoder-based neural ODE strategy been used to develop accelerated surrogate models in a variety of physical applications, including chaotic fluid flows [53], advection-dominated PDEs [54, 55], and stiff chemical kinetics [56, 57]. Moreover, neural ODE surrogates can be effectively combined with neural controllers to tackle complex control problems, as demonstrated in recent works on medical digital twins and optimal control of dynamical systems [58, 59, 60].
@@
Ultimately, previous work has shown how the combination of autoencoders with neural ODEs can be used to generate highly accelerated reduced-order models of physical systems. Despite this, the source of acceleration in the overall modeling strategy remains unclear: the goal of achieving model accuracy from both forecasting and autoencoding perspectives often overshadows the need to identify the contribution of each of these components to the empirically observed model acceleration. Insights into the sources of acceleration provided by the neural ODE based ROM can lead to valuable physical and model-oriented insights, and is the scope of the current study. Recent work has observed the effect of smoothed latent trajectories for advection-dominated systems produced by neural ODE simulations, pointing to a relationship between model acceleration and instrinsic timescale elimination [55]. Similar trends have been shown for neural ODE based surrogate models for stiff chemical kinetics [56]. As such, a more rigorous and quantitative analysis of timescale elimination produced by both autoencoder and neural ODE components is warranted. Additionally, the effect of critical neural ODE training parameters – such as the overall integration time used to evaluate model errors – on the accuracy and degree of timescale elimination produced in the latent space has been largely unexplored in the literature. Lastly, although previous work has demonstrated application of neural ODE based surrogate models for simplified advection-dominated PDEs (e.g., Burger’s equation), extension of this strategy to more complex shock-containing flows remains sparsely explored. To this end, the main contributions of this work are as follows:
@@
    Extend the proposed framework to a challenging and highly advection-dominated test case, specifically 1D detonation wave propagation (considering stiff chemical kinetics), and to a real-world 2D atmospheric dataset, to validate and affirm the observed trends.
It is emphasized that while the approach of combining autoencoders for spatial compression with neural ODEs for learning latent dynamics has been previously explored, this work leverages the general autoencoder-based neural ODE strategy to facilitate a latent time-scale analysis through utilization of dynamical system Jacobians. Specifically, we investigate how the autoencoder architecture, latent space dimensionality, and, most critically, the training trajectory length ($n_{t}$)–parameters often selected heuristically–impact the accuracy and computational speedups achievable in the latent space. Through three distinct test cases, it is demonstrated how $n_{t}$ has a notable influence on the smoothness of latent space trajectories, enabling the use of larger time-steps in the latent space.
@@ 2.1 Surrogate Modeling Task
The ODE in Eq. 2 is interpreted here as a ”ground-truth” representation of the continuous PDE counterpart, meaning the grid is assumed to be resolved enough to properly capture the contribution of all spatiotemporal scales in the physical operators, resulting in a high-dimensional state vector ${\bf u}$ that corresponds to a physical space representation of the state. The motivation for surrogate modeling is that the ground-truth ODE in Eq. 2 is computationally expensive to solve for realistic applications that require fast (near real-time) flow field predictions, such as full-scale design optimization [61, 62, 63] or model-predictive control [64, 65].
@@
It should be noted that accelerated evaluations over the full system using a combined autoencoder and neural ODE strategy can be achieved through both a reduction in system dimensionality through the latent variables (the surrogate ODE operates on a lower-dimensional representation, which facilitates faster evaluations of instantaneous rates), and also an increase in minimum system timescales in the latent space inherent to the functional form of ${\bf G}$ [56] (larger limiting timescales in ${\bf G}$ imply elimination of prohibitive timescales in ${\bf F}$, which in turn allows for larger time-steps to be utilized in the simulation procedure). More specifically, previous work has empirically observed how employing a convolutional autoencoder yields smooth trajectories in the latent space for chaotic PDE-based systems [55]. To this end, within the scope of data-based surrogate modeling, the primary emphasis of this work is to (a) provide a rigorous quantitative analysis on how timescale elimination is achieved in the latent space, (b) identify the role played by key neural ODE training hyper-parameters on this timescale elimination, and (c) understand the manner in which coupling between autoencoder and neural ODE components contributes to forecasting accuracy and latent timescale sensitivity.